OpenStudy (anonymous):

[3.06] If h(x) = −3x − 10, find h(−3). −19 −1 1 2.3

OpenStudy (anonymous):

@AkashdeepDeb

OpenStudy (anonymous):

Same thing, substitute x = -3 into the equation

OpenStudy (anonymous):

-19?

OpenStudy (anonymous):

h(x) = −3x − 10 puttinn x=-3 we have h(-3) = −3(-3) − 10 =9-10=-1

OpenStudy (anonymous):

Remember that - - = +

OpenStudy (anonymous):

oh so 1?

OpenStudy (anonymous):

-1

OpenStudy (anonymous):

Yes